Consider a project lasting one year only. The initial outlay is $1,000 and the expected inflow is $1,210. The opportunity cost of capital is r = 0.21. The borrowing rate is rD = 0.11, and the tax shield per dollar of interest is Tc = 0.21. ( Do not round intermediate calculations. Round your answers to 2 decimal places. Leave no cells blank - be certain to enter "0" wherever required.)

a.
What is the project’s base-case NPV?

b. What is its APV if the firm borrows 31% of the project’s required investment?
Answer #1

I have answered the question below

Please upvote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Solution- A:

Base-case NPV = -1,000 + 1210/1.21 = -1000 + 1000 = 0 ( where 1.21 = 1+cost of capital)

Solution- B:

Interest tax shield = Borrowed amount x Borrowing rate x Tax rate

= 0.31 * 1000 * 0.11 * .21 = $7.161

Interest tax shield Present values of the interest tax shield = (1+ Cost of capital)

= 7.161/(1+ 0.11) = $6.45

APV = Base NPV + PV of tax shield = 0 + 6.45 = $6.45